Excel, please explain the formulas used in excel.Red ID: Student Name: REMBRANT PAINT COMPANY Income Statement For the year ended December 31, 2021 ($ in thousands, except pe
Answer #1


D E F F G H Income statement items year ended 12/31/2021 ($ in thousands) 18,000 10,500 7,500 700 1,200 600 5000 A B с 1 2 3

D E F H Income statement items year ended 12/31/2021 ($ in thousands) =B10 =B13 =H5-H6 =B14 =B15 =B16 =H7-SUM(G9:G11) A B с 1